Nerissa Oberlander is currently enrolled as an undergraduate student in the Computer Science Department at the University of Nevada.

She will receive her B.S. degree in Computer Science from the University of Nevada, Reno in May of 2000.

Besides the paper she will be presenting today, she has a paper that was recently presented at APCOM ’99 (The Applications of Computers in the Mining Industry) and a journal publication that will appear in The International Journal of Surface Mining, Reclamation, and the Environment.

Her research interests include virtual reality, graphics, and simulation.